Circular No. 6641

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bz IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6418, 6639) reports the discovery of an apparent supernova (V ∼ 17.5) on V and R CCD images taken on Apr. 27 by S. Chan on the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997bz is located near Abell 1238 at α = 11^{h}22^{m}25^{s}.46, δ = +1^{o}11′21′′.5 (equinox 2000.0). No star was detected at this position on frames taken on Apr. 11. SN 1997bz is also visible on CCD images taken with the Mount Stromlo 1.27-m telescope on Apr. 30. A nearby star has position end figures 22^{s}.56, 13′24′′.8.

S. Jha, P. Challis, P. Garnavich, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, write: “Spectrograms were obtained by P. Berlind at the 1.5-m Tillinghast reflector on Apr. 30.18 UT and by P. Challis at the Multiple Mirror Telescope on May 2.15. The spectra are heavily contaminated by the host galaxy, but they indicate that SN 1997bz is of type Ia, slightly more than one week after maximum. The redshift derived from Hα emission in the host galaxy is 0.03, implying that it is in the foreground of the cluster (which is at redshift 0.072).”

COMET C/1995 O1 (HALE-BOPP) R. Meier, University of Hawaii (UH); H. Matthews, Joint Astronomy

Centre, Hilo, and National Research Council of Canada; T. Owen and D. Jewitt, UH; M. Senay, University of Massachusetts; and N. Biver, D. Gautier, D. Bockelée-Morvan, and J. Crovisier, Observatoire de Paris, Meudon, report the detection of DCN in comet C/1995 O1 with the James Clerk Maxwell Telescope at Mauna Kea on Apr. 27.9 UT: “The 10σ detection of the DCN 5–4 transition at 362.046 GHz was taken together with a quasisimultaneous measurement of the HCN 4–3 line. Prior observations of HCN and H_{13}CN 4–3 lines will allow an accurate, essentially model-independent determination of the D/H ratio in cometary HCN. Our preliminary analysis leads to an approximate DCN/HCN ratio on the order of 10^{−3}. At the time of the observations, the comet was passing in front of the Taurus dark clouds, and we have been careful to exclude the possibility of contamination by background sources. The detection of DCN is the first detection of cometary deuterium in a molecule other than water. The availability of a D/H ratio in both water (IAUC 6615) and a carbon-nitrogen-bearing molecule may provide fundamental information on the origin and evolution of comets.”

1997 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green